# Stability of quasicrystalline ultracold fermions to dipolar interactions

### v2

@@ I.2 Energies
The unit of energy is defined in terms of the recoil energy of the primary lattice, i.e. $E_{r}\equiv\frac{\hbar^{2}k_{p}^{2}}{2m}\approx 1.1679\times 10^{-29}$ J with $m\approx 39.9640$ Da the mass of 40K atoms. More specifically, we define the unit of energy as $\bar{E}\equiv\frac{\hbar^{2}}{m}\frac{1}{\bar{L}^{2}}=\frac{2E_{r}}{\pi^{2}}\approx 0.237\times 10^{-19}$ J. In typical experiments with quasiperiodic optical lattices, the depths are varied in regimes of up to around 8 recoil energies for the primary lattice, and around 1 recoil energy for the detuning lattice. In our simulations, we probe similar regimes: $V_{p}\in[15\bar{E},60\bar{E}]\approx[3.0E_{r},12.2E_{r}]$, $V_{d}\in[0\bar{E},6\bar{E}]=[0E_{r},1.2E_{r}]$. When we turn on dipolar interactions, we probe the regimes with strength $W\in[0\bar{E},10.0\bar{E}]\approx[0E_{r},2E_{r}]$, which is in accordance to the values that can be achieved in near-term ultracold dipolar quantum simulators.
@@ I.3 Time
The unit of time is also defined from the unit of length as $\bar{t}\equiv\frac{m\hat{L}^{2}}{\hbar}=\frac{m\lambda_{p}^{2}}{4\hbar}=4.46\times 10^{-5}$ s. In our simulations, we run time evolutions up until around $t\approx 250\bar{t}\approx 11.2$ ms. Note that this is much shorter compared to experimental runs (in the order of a couple of seconds), but it allows us to probe all the essential features of the system because its dynamics is quite fast. In fact, the core features of the phase diagram can be faithfully extracted already at $t=50\bar{t}\approx 2.24$ ms. Furthermore, since are dealing with a smaller system size compared to the experiments, we need to extract expansion measurements at a much shorter time scale, before the particles hit the boundaries of the system.
@@ I.4 Simulation time
4 6 32.692
4 8 13.157
6 8 13.871
6 10 6.958
Table 1: Average final simulation times for the dynamics of $N$ fermions and $M$ orbitals in a quasiperiodic optical lattice obtained on 2.4 GHz processors for runs of 5 days.
@@ III Orbital convergence
where the eigenvectors $\phi_{i}^{\mathrm{(NO)}}(x)$ are termed natural orbitals. The occupations are ranked from most occupied to least occupied, and are normalized such that they sum up to one. A measure of the exactness of the MCTDH-X ansatz is given by orbital convergence: an MCTDH-X calculation with $M$ orbitals can be declared converged in orbitals number within a given threshold $t$, if the population of an additional orbital stays within that threshold, i.e. $\rho_{M+1}<t$. Typical threshold values are $t\gg\frac{1}{M}$, e.g. for $M=9$, $t=0.01$.
@@
Figure S3: Time evolution of orbital occupations in each phase for increasing interaction strength $W$: extended phase (left column), localized phase (middle column), and intermediate phase (right column). The interaction strength is (a)-(c) $W=0.02$$E_{r}$, (d)-(f) $W=0.2$$E_{r}$, (g)-(i) $W=2.0$$E_{r}$. The potential depths are (a), (d), (g): $V_{p}=3.0$$E_{r}$, $V_{d}=0.0$$E_{r}$, (b), (e), (h): $V_{p}=8.1$$E_{r}$, $V_{d}=1.2$$E_{r}$, (c), (f), (i): $V_{p}=6.1$$E_{r}$, $V_{d}=0.4$$E_{r}$. The symbols indicate the phase of each column and where in parameter space they are taken from (cf. Fig. S1): extended phase (circle), intermediate phase (star), and localized phase (square).
IV Results for different particle number and orbitals
@@
Figure S4: Phase diagrams for $N$ ultracold fermions in a 1D quasiperiodic potential. (a) $N=4$ particles, $M=6$ orbitals. (b) $N=4$ particles, $M=8$ orbitals. (c) $N=6$ particles, $M=8$ orbitals. (d) $N=6$ particles, $M=10$ orbitals. For the localized phase, the time-averaged imbalance parameter $\mathcal{I}\right>$ is further plotted as a continuous color scheme.
@@
V Density expansion dynamics
@@
In this section, we present direct results for the density expansion dynamics that complement the indirect expansion measure obtained from the calculation of the edge density and illustrated in the main text. In Fig. S5, we show the behavior of the density expansion in time $\mathcal{E}(t)$ for various points in parameter space belonging to different phases. The expansion is calculated as
@@
From the figure S5(a), we can distinguish three different behaviors for the data points analyzed in the main text and corresponding to the three different phases (extended, intermediate, and localized). The extended phase shows a very rapid and linear expansion. The localized phase, on the contrary, shows practically zero expansion. The intermediate phase (highlighted by the grey rectangle) shows instead a slow, sublinear expansion that becomes progressively more chaotic as dipolar interactions increase. Upon closer inspection in S5(b), the intermediate phase shows a slight slowing down of expansion when interactions are increased. This is a sign of the increasing interference of the tail of the dipole-dipole interactions.
Nevertheless, locating the phase boundary between extended and intermediate phase is not so easy based on the expansion dynamics alone. A pictorial demonstration of this is offered in S5(c), where for $W=0.2E_{r}$ we plot with higher resolution the expansion dynamics of the density for parameters across the transition from extended to localized phase (shown in the inset). Indeed, while we progressively move towards the intermediate phase, the monotonic and rapid density expansion continuously slows down, and starts to acquire a non-monotonic character. These results point towards a crossover from extended to intermediate phase, rather than a sharp transition. Indeed, the hallmark of the intermediate phase is rather a coexistence of both extended and localized states. In the long-time limit, the extended states will always expand indefinitely. However, localized states will remain localized in the initial configuration. This is where the measurement of the imbalance comes into play to determine which regions of parameter space belong to the intermediate phase.
@@
VI Behavior of average kinetic and interaction energy
@@
VII Correlation dynamics
In this section we present additional information about the dynamics of correlations in each phase and for increasing dipolar interaction strengths. These figures complement the ones in the main text by providing information at earlier and later times ($t=0.089$ ms and $t=2.23$ ms). The 2-RDM for the noninteracting case ($W=0$) and for three values of the interactions ($W=0.01$$E_{r}^{p}$, $W=0.10$$E_{r}^{p}$, and $W=1.01$$E_{r}^{p}$) is depicted in Figs. S7, S8, S9, and S10, respectively. All calculations were performed with $M=9$ orbitals.

### v3

@@ I.2 Energies
The unit of energy is defined in terms of the recoil energy of the primary lattice, i.e. $E_{r}\equiv\frac{\hbar^{2}k_{p}^{2}}{2m}\approx 2.89\times 10^{-30}$ J with $m\approx 2.673$ 10-25 kg the mass of 161Dy atoms. The corresponding recoil frequency is $\nu_{r}=E_{r}/h\approx 4.36kHz$. More specifically, we define the unit of energy as $\bar{E}\equiv\frac{\hbar^{2}}{m}\frac{1}{\bar{L}^{2}}=\frac{2E_{r}}{\pi^{2}}\approx 5.86\times 10^{-31}$ J, or in frequency, $\nu_{\bar{E}}=884$ Hz.
In typical experiments with quasiperiodic optical lattices, the depths are varied in regimes of up to around 8 recoil energies for the primary lattice, and around 1 recoil energy for the detuning lattice. In our simulations, we probe similar regimes: $V_{p}\in[15\bar{E},60\bar{E}]\approx[3.0E_{r},12.2E_{r}]$, $V_{d}\in[0\bar{E},6\bar{E}]=[0E_{r},1.2E_{r}]$. In terms of frequencies, $V_{p}\in[13.1,53.2]$ kHz, $V_{d}\in[0,5.2]$ kHz.
When we turn on dipolar interactions, we probe the regimes with strength $W\in[0\bar{E},10.0\bar{E}]\approx[0E_{r},2E_{r}]$ (in terms of frequencies: $W\in[0,8.7]kHz$), which is in accordance to the values that can be achieved in near-term ultracold dipolar quantum simulators.
@@ I.3 Time
[…]
@@ I.4 Simulation time
4 6 130.8
4 8 52.6
6 8 55.5
6 10 27.8
Table 1: Average final simulation times for the dynamics of NN fermions and $M$ orbitals in a quasiperiodic optical lattice obtained on 2.4 GHz processors for runs of 5 days.
@@ III Orbital convergence
where the eigenvectors $\phi_{i}^{\mathrm{(NO)}}(x)$ are termed natural orbitals. The occupations are ranked from most occupied to least occupied, and are normalized such that they sum up to one. A measure of the exactness of the MCTDH-X ansatz is given by orbital convergence: an MCTDH-X calculation with $M$ orbitals can be declared converged in orbitals number within a given threshold $C$, if the population of an additional orbital stays within that threshold, i.e. $\rho_{M+1}<t$. Typical threshold values are $C\gg\frac{1}{M}$, e.g. for $M=9$, $t=0.01$.
@@
[…]
IV Orbital profiles
In this section, we visualize the profiles of the $M=9$ orbitals used to calculate the dynamics of $N=4$ particles presented in the main text. This visualization should help the reader understand that the orbitals are truly optimized both in space and in time, and might end up acquiring a very nonlocal profile when the optical lattices are shallow and the lattice description is a very coarse approximation.
Exemplary orbitals are depicted in Fig. S4. Panels (a)-(b) visualize the orbitals at time $t=0.0$ (their profile is identical for all phases since they all start from the same initial state in the time evolution) – obtained from imaginary time evolution into the ground state corresponding to four sites in the center of the optical lattice. Already here we can appreciate that the orbitals are delocalized over many sites. This freedom of delocalization is what gives MCTDH orbital a high degree of expressivity. Panels (c)-(d) show the orbitals at time $t=50\bar{t}$ for a data point in the extended phase, while panel (e)-(f) shows the orbitals at time $t=50\bar{t}$ for a data point in the localized phase (note the different scale in the $x$-axis). As we can see from the plots, in the extended phase the orbitals become highly nonlocal during time evolution, to follow the density expansion. In the localized phase, instead, the orbitals retain most of their weight in the center of the system.
Figure S4: Examples of real (left columns) and imaginary (right columns) amplitudes for the MCTDH orbitals used in the many-body expansion. The index $i$ labels the orbitals from largest to smallest contribution. (a)-(b) Orbital profiles for the initial state at $t=0$. (c)-(d) Orbital profiles at time $t=50\bar{t}$ for the system in the extended phase with $V_{p}=3.0E_{r}$, $V_{d}=0.4E_{r}$, $W=0.02E_{r}$. (e)-(f) Orbital profiles at time $t=50\bar{t}$ for the system in the localized phase with $V_{p}=8.0E_{r}$, $V_{d}=1.2E_{r}$, $W=0.02E_{r}$.
V Results for different particle and orbital number
@@
Figure S5: Phase diagrams for $N$ ultracold fermions in a 1D quasiperiodic potential. (a) $N=4$ particles, $M=6$ orbitals. (b) $N=4$ particles, $M=8$ orbitals. (c) $N=6$ particles, $M=8$ orbitals. (d) $N=6$ particles, $M=10$ orbitals. For the localized phase, the time-averaged imbalance parameter $\left$$\mathcal{I}\right$> is further plotted as a continuous color scheme.
@@
VI Effective nearest-neighbor hopping
In this section, we estimate the nearest-neighbor hopping amplitude $J_{\mathrm{NN}}$ in the lattice regime we simulate using a semiclassical approximation based on localized Wannier functions. Although our simulations are performed in the continuum and the primary optical lattice is rather shallow (typically $V_{p}\in[3E_{r},8E_{r}]$), it is still instructive to compare with the tight-binding regime for orientation.
In deep lattice limits, the lowest-band Wannier functions can be approximated by Gaussians, and the hopping between adjacent sites can be estimated analytically. We follow the standard approximation (see e.g. Morsch and Oberthaler, Rev. Mod. Phys. 78, 179 (2006)):
$J_{\mathrm{NN}}\approx\frac{4}{\sqrt{\pi}}E_{r}\left(\frac{V_{p}}{E_{r}}\right)^{3/4}\exp\left(-2\sqrt{\frac{V_{p}}{E_{r}}}\right).$ (S4)
Here, $V_{p}$ is the depth of the primary optical lattice, and $E_{r}$ is the recoil energy defined in the sections above.
To illustrate the scaling, let us consider a few representative lattice depths:
  • For $V_{p}=3E_{r}$:
    $J_{\mathrm{NN}}\approx 0.142E_{r}\Rightarrow J_{\mathrm{NN}}/h\approx 590\,\mathrm{Hz}$,  $\hbar/J_{\mathrm{NN}}\approx 1.69\,\mathrm{ms}$
  • For $V_{p}=6E_{r}$:
    $J_{\mathrm{NN}}\approx 0.0645E_{r}\Rightarrow J_{\mathrm{NN}}/h\approx 268\,\mathrm{Hz}$,  $\hbar/J_{\mathrm{NN}}\approx 3.73\,\mathrm{ms}$
  • For $V_{p}=9E_{r}$:
    $J_{\mathrm{NN}}\approx 0.030E_{r}\Rightarrow J_{\mathrm{NN}}/h\approx 125\,\mathrm{Hz}$,  $\hbar/J_{\mathrm{NN}}\approx 8.00\,\mathrm{ms}$
These results show that the hopping times remain much smaller than the maximum time scale that we probe in our numerics (up to 45 ms).
Furthermore, we compare these hopping energies with the strength of dipolar interactions used in our simulations. For example, in the strongly interacting case $V_{d}=2E_{r}$, we find:
$\frac{V_{d}}{J_{\mathrm{NN}}}\approx\frac{2.0}{0.0645}\approx 31.$
Thus, the dipolar interaction energy scale dominates over the kinetic energy scale set by $J_{\mathrm{NN}}$, confirming that we operate deep in the strongly interacting regime. This supports our conclusion that the behavior observed in our simulations – especially in the localized regime – cannot be captured perturbatively from a weakly interacting tight-binding perspective and continuum formulations like MCTDH-X become necessary.
VII Dynamics of observables
In this section, we visualize the full time evolution of the three main observables mentioned in the main text: the imbalance $\mathcal{I}$, the edge density $\mathcal{D}$, and the expansion $\mathcal{E}$.
Each panel of Fig. S6 shows expansion dynamics for three or four representative parameter choices in the extended, intermediate, and localized phase, respectively, and for a different value of dipolar interaction strength: (a) $W=0.0E_{r}$, (b) $W=0.02E_{r}$, (c) $W=0.2E_{r}$, (d) $W=2.0E_{r}$. To plot all curves in the same panels, and since the expansion is quite rapid in the extended phase, we only show times up to $t=2.2ms$. We can clearly distinguish the three qualitatively different behaviors captured by the three phases presented in the main text. In the extended phase, the density rapidly expands to reach the system boundaries. In the localized phase, no expansion is observed (we have verified that the expansion remains negligible for longer times). In the intermediate phase, expansion occurs but it is much slower than in the extended phase, and does not reach the system boundaries for the time scales probed by our numerics.
Figure S6: Density expansion dynamics for three different parameter values corresponding to the extended phase (purple curves), the intermediate phase (cyan/blue curves) and the localized phases (olive curve: DSBS state, red curve: standard localized phase). The different panels depict the dynamics for different values of dipolar interactions: (a) $W=0.0E_{r}$, (b) $W=0.02E_{r}$, (c) $W=0.2E_{r}$, (d) $W=2.0E_{r}$. The dashed lines are the raw data, while the thick solid lines are running averages.
Fig. S7 depicts the behavior of the edge density, another expansion proxy which provides a numerically more stable quantity to track the amount of expansion (and allows us to plot different phases on the same scale for longer times). Each panel shows three or four representative parameter choices in the extended, intermediate, and localized phase, respectively, and for a different value of dipolar interaction strength: (a) $W=0.0E_{r}$, (b) $W=0.02E_{r}$, (c) $W=0.2E_{r}$, (d) $W=2.0E_{r}$. The behavior of the three different phases is similar to that of the raw expansion. iI the extended phase, the edge density increases fast to a saturation value. In the localized phase, the edge density stays essentially at zero through time evolution. In the intermediate phase, the edge density tends to increase, but a slower rate than in the extended phase. Note also how stronger interactions tend to inhibit the growth rate of the edge density, consistent with the localizing role of strong dipolar repulsions.
Figure S7: Edge density dynamics for three different parameter values corresponding to the extended phase (purple curves), the intermediate phase (cyan/blue curves) and the localized phases (olive curve: DSBS state, red curve: standard localized phase). The different panels depict the dynamics for different values of dipolar interactions: (a) $W=0.0E_{r}$, (b) $W=0.02E_{r}$, (c) $W=0.2E_{r}$, (d) $W=2.0E_{r}$. The dashed lines are the raw data, while the thick solid lines are running averages.
Fig. S8 shows the behavior of the density imbalance, which measures the wave function localization. Each panel shows three or four representative parameter choices in the extended, intermediate, and localized phase, respectively, and for a different value of dipolar interaction strength: (a) $W=0.0E_{r}$, (b) $W=0.02E_{r}$, (c) $W=0.2E_{r}$, (d) $W=2.0E_{r}$. The imbalance is an inherently noisier quantity because of the ambiguity of defining lattice sites in a continuum optical lattice. To compensate for that, in selecting the value used to classify the different phases, we typically consider averages over the last 50 time steps. The averages nicely recollect three distinct cases. iI the extended phase, the initial imbalance rapidly decays towards zero and then oscillates close to that value for all subsequent times. In the localized phase, the initial imbalance is mostly retained with a few oscillations. In particular, strong interactions preserve much better the initial imbalance [cf. Fig. S8(d)]. In the intermediate phase, the imbalance exhibits a larger reduction from its initial value, but remains well above zero throughout the dynamics.
Figure S8: Imbalance dynamics for three different parameter values corresponding to the extended phase (purple curves), the intermediate phase (cyan/blue curves) and the localized phases (olive curve: DSBS state, red curve: standard localized phase). The different panels depict the dynamics for different values of dipolar interactions: (a) $W=0.0E_{r}$, (b) $W=0.02E_{r}$, (c) $W=0.2E_{r}$, (d) $W=2.0E_{r}$. The dashed lines are the raw data, while the thick solid lines are running averages.
VIII More details on the density expansion dynamics
@@
In this section, we present additional results for the density expansion dynamics that complement the indirect expansion measure obtained from the calculation of the edge density and illustrated in the main text. In Fig. S9, we show the behavior of the density expansion in time $\mathcal{E}(t)$ for various points in parameter space belonging to different phases. The expansion is calculated as
@@
From the figure S9(a), we can distinguish three different behaviors for the data points analyzed in the main text and corresponding to the three different phases (extended, intermediate, and localized). The extended phase shows a very rapid and linear expansion. The localized phase, on the contrary, shows practically zero expansion. The intermediate phase (highlighted by the grey rectangle) shows instead a slow, sublinear expansion that becomes progressively more chaotic as dipolar interactions increase. Upon closer inspection in S9(b), the intermediate phase shows a slight slowing down of expansion when interactions are increased. This is a sign of the increasing interference of the tail of the dipole-dipole interactions.
Nevertheless, locating the phase boundary between extended and intermediate phase is not so easy based on the expansion dynamics alone. A pictorial demonstration of this is offered in S9(c), where for $W=0.2E_{r}$ we plot with higher resolution the expansion dynamics of the density for parameters across the transition from extended to localized phase (shown in the inset). Indeed, while we progressively move towards the intermediate phase, the monotonic and rapid density expansion continuously slows down, and starts to acquire a non-monotonic character. These results point towards a crossover from extended to intermediate phase, rather than a sharp transition. Indeed, the hallmark of the intermediate phase is rather a coexistence of both extended and localized states. In the long-time limit, the extended states will always expand indefinitely. However, localized states will remain localized in the initial configuration. This is where the measurement of the imbalance comes into play to determine which regions of parameter space belong to the intermediate phase.
@@
IX Behavior of average kinetic and interaction energy
@@
X Correlation dynamics
[…]